OpenStudy (anonymous):

Solve for x: -2(x - 3) = 10 x = -2 x = 2 x = -8 x = 8

OpenStudy (anonymous):

First expand the brackets, then solve for x.

OpenStudy (ash2326):

@kenbaby We have \[-2(x-3)=10\] Divide both sides by -2 \[-2(x-3)/(-2)=10/(-2)\] We get \[x-3=-5\] Add 3 to both the sides \[x-3+3=-5+3\] Can you proceed from here?

OpenStudy (anonymous):

so would it be the first one?